A 50 kg block of wood rest on the top of the smooth plane whose length is 3m and 

whose altitude is 0.8m. How long will it take for the block to slide to the bottom of 

the plane when release?


1
Expert's answer
2021-03-10T17:19:31-0500

Length of incline(s) = 3 m

Vertical rise of incline = 0.8 m

Angle of inclination = "\theta=sin^{-1}{\{\dfrac{Vertical \space Rise}{Length}\}}=15.466\degree"

Acceleration of the block along the incline(a) = "g\sin\theta"

From newton's laws of motion

"s=ut+\dfrac{1}{2}at^2"

"\Rightarrow t=\sqrt{\dfrac{2s}{a}}"

"\Rightarrow t=\sqrt{\dfrac{6}{2.6133}}=1.51s"